# Bondi-type accretion onto a Kerr black hole in the kinetic regime

### v1

@@ Abstract
We derive an exact solution representing a Bondi-type stationary accretion of a kinetic (Vlasov) gas onto the Kerr black hole. The solution is exact in the sense that relevant physical quantities, such as the particle current density or the accretion rates, are expressed as explicit integrals, which can be evaluated numerically. Whereas the angular momentum accretion rate is shown to vanish exactly, we provide an analytic approximation which allows us to obtain simple formulas for the mass and energy accretion rates. These formulas are used to derive characteristic time scales of the black hole mass growth and the associated spin-down in two different scenarios: assuming that the ambient energy density is either constant or decreases on a cosmological scale.
@@ II Kinetic description of matter in the Kerr spacetime
where $\rho^{2}:=r^{2}+a^{2}\cos^{2}\vartheta$, and $M$ and $a$ denote the black hole mass and the rotation parameter, respectively. We restrict ourselves to the sub-extremal case, such that the spin parameter $\alpha:=a/M$ lies in the inverval $1<\alpha<1$. The Kerr spacetime admits the two Killing vector fields $k=\partial_{t}$ and $\eta=\partial_{\varphi}$.
@@
Figure 1: Compression ratio at the meridional plane close to the black hole horizon (white dashed line) for the Maxwell-Jüttner model with $z=1$. We denote $x=r/M$. The solid dark lines mark contours of constant $n$. Left: $\alpha=0.5$; right: $\alpha=0.95$. This plot can be compared with Fig. 11 in [20] obtained for monoenergetic distributions.
@@ III Bondi-type solution
Figure 1 shows the compression ratio $n/n_{\infty}$ in the meridional plane for black holes with spin parameters $\alpha=0.5$ and $\alpha=0.95$ and the inverse temperature $z=1$. A polar-to-equatorial asymmetry is clearly visible in the second case, the compression ratio at the equator being higher than at the poles for points lying close to the horizon.
@@ IV Accretion rates
$\mathcal{\dot{N}}=-\pi\int\limits_{m}^{\infty}dEF_{\infty}(E)\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi(Q_{c})^{2}.$ (17)
@@
and
$\mathcal{\dot{J}}$ $=$ $-\frac{2\pi}{3}\int\limits_{m}^{\infty}dEF_{\infty}(E)\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi\cos\chi(Q_{c})^{3}$ (19)
$=$ $0.$
The vanishing of the angular momentum accretion rate $\dot{\mathcal{J}}$ is due to the fact that $Q_{c}$ is a smooth function of $\sin\chi$.
@@
$\frac{\dot{\mathcal{N}}}{n_{\infty}}$ $=$ $-\frac{\pi zM^{2}}{K_{2}(z)}\int_{1}^{\infty}\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}\right]e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{3}),$ (20a)
$\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}$ $=$ $-\frac{\pi z^{2}M^{2}}{zK_{1}(z)+3K_{2}(z)}\int_{1}^{\infty}\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}\right]\varepsilon e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{3}),$ (20b)
@@
An even simpler expression can be obtained in the low ($z\to\infty$) and high ($z\to 0$) asymptotic temperature limits. Neglecting terms that are cubic or higher order in $\alpha$, one gets
@@
The exact values of the energy accretion rates are illustrated in Fig. 2, both for small and large values of $z$. They depend weakly on the black hole spin parameter and hence remain in excellent agreement with the results of the slow-rotation approximation, beyond the expected range of $\alpha$. We checked that for several values of $z$ in the range $z\in(0,\infty)$, the difference between the results obtained within the slow-rotation approximation and the exact numerical values is smaller than $4\%$ for black holes with $\alpha\leq 0.99$.
@@ V Discussion
$\frac{dM}{dt}=-\dot{\mathcal{E}},\qquad\frac{dJ}{dt}=-\dot{\mathcal{J}}=0.$ (23)
Consequently, the black hole angular momentum remains constant: $J=\alpha M^{2}=\alpha_{0}M_{0}^{2}$, where $\alpha_{0}$ and $M_{0}$ denote, respectively, the spin parameter and the black hole mass at the initial time $t_{0}$. Assuming $\dot{\mathcal{E}}=-\varepsilon_{\infty}\mathcal{A}(1-\mathcal{B}\alpha^{2})M^{2}$ as in Eqs. (22) with $\mathcal{A}=16\sqrt{2\pi z}$, $\mathcal{B}=1/16$ for $z\to\infty$ and $\mathcal{A}=27\pi$, $\mathcal{B}=2/27$ for $z\to 0$, one obtains (restoring SI units)
$\frac{dM}{dt}=\frac{G^{2}\mathcal{A}}{c^{5}}\left(M^{2}-\mathcal{B}\alpha_{0}^{2}\frac{M_{0}^{4}}{M^{2}}\right)\varepsilon_{\infty}.$ (24)
We conclude this letter with two applications of Eq. (24). The first one considers a cosmological scenario consisting of the growth of primordial black holes, due to the accretion of hot dark matter particles, following [25] and [29]. In this case, $\mathcal{A}=27\pi$ is constant and $\varepsilon_{\infty}=3c^{2}/(32\pi Gt^{2})$, and integrating Eq. (24) from $t_{0}$ to $t_{X}$, corresponding to a mass growth from $M_{0}$ to $M=XM_{0}$ and a decrease of the black hole spin parameter from $\alpha_{0}$ to $\alpha=\alpha_{0}/X^{2}$, yields
$\mathcal{F}_{X}(\alpha_{0})=\frac{81\kappa}{32}\left(1-\frac{t_{0}}{t_{X}}\right).$ (25)
Here $\kappa$ is a dimensionless parameter of the order of unity defined by $M_{0}=\kappa c^{3}t_{0}/G$, as in [29], and
$\mathcal{F}_{X}(\alpha_{0})$ $=$ $\int_{1}^{X}\frac{dy}{y^{2}(1-\mathcal{B}\alpha_{0}^{2}/y^{4})}$ (26)
$=$ $-\frac{1}{2\mathcal{C}}\left[\arctan\left(\frac{\mathcal{C}}{y}\right)+\mbox{arctanh}\left(\frac{\mathcal{C}}{y}\right)\right]_{y=1}^{y=X},\qquad$
with $\mathcal{C}=(\mathcal{B}\alpha_{0}^{2})^{1/4}$ (note that $\mathcal{C}<1$ and $X\geq 1$, so that this expression is well defined). Assuming that $81\kappa/(32\mathcal{F}_{\infty}(\alpha_{0}))>1$, the mass diverges and the black hole spin parameter $\alpha$ tends to zero in a finite time $t_{\infty}$, given by
$\frac{t_{0}}{t_{\infty}}=1-\frac{32\mathcal{F}_{\infty}(\alpha_{0})}{81\kappa},$ (27)
where
$\mathcal{F}_{\infty}(\alpha_{0})$ $=$ $\frac{\mathrm{arctan}(\mathcal{C})+\mathrm{arctanh}(\mathcal{C})}{2\mathcal{C}}$ (28)
$=$ $1+\frac{\mathcal{B}}{5}\alpha_{0}^{2}+\mathcal{O}(\mathcal{B}^{2}\alpha_{0}^{4}),$
and $\mathcal{B}=2/27$. Since $\mathcal{F}_{\infty}(\alpha_{0})$ lies very close to one ($\mathcal{F}_{\infty}(\alpha_{0}=1)\approx 1.01546$), one is led to the same conclusion as in the Schwarzschild model: if $\kappa\sim 1$ the accretion process by primordial black holes is robust, even if they are highly spinning initially.
The second scenario corresponds to a supermassive black hole which accretes dark matter from its environment. Assuming a reservoir of constant density and low temperature and integrating Eq. (24), one finds that the time required for the black hole mass to grow by a factor $X$ (and the black hole spin parameter to decrease from $\alpha_{0}$ to $\alpha=\alpha_{0}/X^{2}$) is $t_{X}=\mathcal{F}_{X}(\alpha_{0})c^{5}/(G^{2}M_{0}\mathcal{A}\varepsilon_{\infty})$, where $\mathcal{F}_{X}(\alpha_{0})$ is defined as in Eq. (26) and where for simplicity we have set $t_{0}=0$. Therefore, the characteristic time scale $t_{\infty}$ for having a significant mass growth is related to the parameter $z$ according to
@@
$\sqrt{z}=\frac{c^{5}\mathcal{F}_{\infty}(\alpha_{0})}{\sqrt{512\pi}G^{2}M_{0}\varepsilon_{\infty}t_{\infty}},$ (29)

### v2

@@ Abstract
We derive an exact solution representing a Bondi-type stationary accretion of a kinetic (Vlasov) gas onto the Kerr black hole. The solution is exact in the sense that relevant physical quantities, such as the particle current density or the accretion rates, are expressed as explicit integrals, which can be evaluated numerically. We provide an analytic approximation which allows us to obtain simple formulas for the mass, energy, and angular momentum accretion rates. These formulas are used to derive characteristic time scales of the black hole mass growth and the associated spin-down in two different scenarios: assuming that the ambient energy density is either constant or decreases on a cosmological scale.
@@ II Kinetic description of matter in the Kerr spacetime
where $\rho^{2}:=r^{2}+a^{2}\cos^{2}\vartheta$, and $M$ and $a$ denote the black hole mass and the rotation parameter, respectively. We restrict ourselves to the sub-extremal case, such that the spin parameter $\alpha:=a/M$ lies in the interval $1<\alpha<1$. The Kerr spacetime admits the two Killing vector fields $k=\partial_{t}$ and $\eta=\partial_{\varphi}$.
@@
Figure 1: Compression ratio at the meridional plane close to the black hole horizon (white dashed line) for the Maxwell-Jüttner model with $z=1$ and $\alpha=0.95$. We denote $x=r/M$. The solid dark lines mark contours of constant $n$. This plot can be compared with Fig. 11 in [20] obtained for monoenergetic distributions.
@@ III Bondi-type solution
Figure 1 shows the compression ratio $n/n_{\infty}$ in the meridional plane for a black hole with spin parameter $\alpha=0.95$ and the inverse temperature $z=1$. A polar-to-equatorial asymmetry is clearly visible, the compression ratio at the equator being higher than at the poles for points lying close to the horizon.
@@ IV Accretion rates
$\mathcal{\dot{N}}$ $=$ $-\pi\int\limits_{m}^{\infty}dEF_{\infty}(E)\int\limits_{0}^{\pi}d\vartheta\sin\vartheta\int\limits_{0}^{2\pi}d\chi(Q_{c})^{2},$ (17a)
@@
$\mathcal{\dot{J}}$ $=$ $-2\pi\int\limits_{m}^{\infty}dEF_{\infty}(E)\int\limits_{0}^{\pi}d\vartheta\sin^{2}\vartheta$ (17c)
$\times\int\limits_{0}^{2\pi}d\chi\left[\frac{1}{3}\sin\chi Q_{c}^{3}+\frac{a}{2}\sin\vartheta EQ_{c}^{2}\right].$
@@
$\frac{\dot{\mathcal{N}}}{n_{\infty}}$ $=$ $-\frac{\pi zM^{2}}{K_{2}(z)}\int_{1}^{\infty}\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}\right]e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{4}),$ (18a)
$\frac{\dot{\mathcal{E}}}{\varepsilon_{\infty}}$ $=$ $-\frac{\pi z^{2}M^{2}}{zK_{1}(z)+3K_{2}(z)}\int_{1}^{\infty}\frac{\left(x_{c}^{(0)}\right)^{2}}{x_{c}^{(0)}-3}\left[1-\frac{2\alpha^{2}}{\left(x_{c}^{(0)}\right)^{2}\left(6-x_{c}^{(0)}\right)}\right]\varepsilon e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{4}),$ (18b)
$\frac{\dot{\mathcal{J}}}{\varepsilon_{\infty}}$ $=$ $\frac{4\pi\alpha z^{2}M^{3}}{3[zK_{1}(z)+3K_{2}(z)]}\int_{1}^{\infty}\left(\frac{x_{c}^{(0)}}{x_{c}^{(0)}-3}\right)^{\frac{3}{2}}\left\{1+\frac{3\alpha^{2}\left(x_{c}^{(0)}-3\right)\left[\left(x_{c}^{(0)}\right)^{2}-24\right]}{5\left(x_{c}^{(0)}-6\right)^{3}\left(x_{c}^{(0)}\right)^{2}}\right\}e^{-z\varepsilon}d\varepsilon+\mathcal{O}(\alpha^{4}),$ (18c)
@@
An even simpler expression can be obtained in the low ($z\to\infty$) and high ($z\to 0$) asymptotic temperature limits. Neglecting terms that are quartic or higher order in $\alpha$, one gets
@@
$\lim\limits_{z\rightarrow\infty}\frac{1}{\sqrt{2\pi z}}\frac{\mathcal{\dot{J}}}{\varepsilon_{\infty}}$ $=$ $\alpha\left(\frac{32}{3}+\frac{2}{5}\alpha^{2}\right)M^{3},$ (20b)
@@
$\lim\limits_{z\rightarrow 0}\frac{\mathcal{\dot{J}}}{\varepsilon_{\infty}}$ $=$ $36\pi\alpha M^{3}.$ (20d)
The exact values of the energy accretion rates are illustrated in Fig. 2, both for small and large values of $z$. They depend weakly on the black hole spin parameter and hence remain in excellent agreement with the results of the slow-rotation approximation, beyond the expected range of $\alpha$. We checked that for several values of $z$ in the range $z\in(0,\infty)$, the difference in the energy and angular momentum accretion rates between the results obtained within the slow-rotation approximation and the exact numerical values is smaller than $4\%$ for the energy accretion rate and below $5\%$ for angular momentum accretion rate for black holes with $\alpha\leq 0.99$.
@@ V Discussion
$\frac{dM}{dt}$ $=$ $-\dot{\mathcal{E}}=\varepsilon_{\infty}\mathcal{A}(1-\mathcal{B}\alpha^{2})M^{2},$ (21a)
$\frac{dJ}{dt}$ $=$ $-\dot{\mathcal{J}}=-2\delta\varepsilon_{\infty}\mathcal{A}(1+\gamma\alpha^{2})M^{3}\alpha,$ (21b)
up to third order in the rotation parameter, where
$\mathcal{A}$ $=$ $16\sqrt{2\pi z},\quad\mathcal{B}=\frac{1}{16},\quad\delta=\frac{1}{3},\quad\gamma=\frac{3}{80},\quad z\to\infty,$
$\mathcal{A}$ $=$ $27\pi,\quad\mathcal{B}=\frac{2}{27},\quad\delta=\frac{2}{3},\quad\gamma=0,\quad z\to 0.$
Equations (21) imply a decrease of the spin parameter magnitude at the rate
@@
$\dot{\alpha}=-2\varepsilon_{\infty}\mathcal{A}(1+\delta)(1-\mathcal{C}\alpha^{2})M\alpha,\quad\mathcal{C}=\frac{\mathcal{B}-\delta\gamma}{1+\delta},$ (21c)
where we have used the relation $J=\alpha M^{2}$. Equations (21a) and (21c) govern the time evolution of the black hole mass and spin parameter, and they yield the following relation:
$\frac{M_{0}}{M}=\left(\frac{\alpha}{\alpha_{0}}\right)^{\frac{1}{2+2\delta}}\left(\frac{1-\mathcal{C}\alpha^{2}}{1-\mathcal{C}\alpha_{0}^{2}}\right)^{\frac{\mathcal{B}-\mathcal{C}}{4(1+\delta)\mathcal{C}}},$ (22)
where $\alpha_{0}$ and $M_{0}$ denote, respectively, the spin parameter and the black hole mass at the initial time $t_{0}$. Substituting the resulting expression for the mass $M$ in Eq. (21c), one obtains an equation for $\alpha$ which, in principle, can be integrated.
We conclude this letter with two applications of Eqs. (21). For these applications, it will turn out to be sufficient to neglect the relatively small term $\mathcal{B}\alpha^{2}$ in Eq. (21a) which decreases during the evolution. Hence, restoring SI units and neglecting terms that are quadratic or higher order in the rotation parameter, we can integrate Eq. (21a) from $t_{0}$ to $t_{X}$, corresponding to a mass growth from $M_{0}$ to $M=XM_{0}$ and a decrease of the spin parameter from $\alpha_{0}$ to $\alpha=\alpha_{0}/X^{2+2\delta}$, as
$1-\frac{1}{X}=\frac{G^{2}}{c^{5}}\mathcal{A}M_{0}\int_{t_{0}}^{t_{X}}\varepsilon_{\infty}(\bar{t})d\bar{t},$ (23)
where we have allowed for the energy density to be time-dependent. The first application considers a cosmological scenario consisting of the growth of primordial black holes, due to the accretion of hot dark matter particles, following [25] and [29]. In this case, $\mathcal{A}=27\pi$ is constant and $\varepsilon_{\infty}=3c^{2}/(32\pi Gt^{2})$. It then follows from Eq. (23) that the mass diverges and the black hole spin parameter $\alpha$ tends to zero in a finite time $t_{\infty}$, given by
$\frac{t_{0}}{t_{\infty}}=1-\frac{32}{81\kappa},$ (24)
where $\kappa$ is a dimensionless parameter of the order of unity defined by $M_{0}=\kappa c^{3}t_{0}/G$, as in [29], and it was assumed that $81\kappa>32$. Therefore, one is led to the same conclusion as in the Schwarzschild model: if $\kappa\sim 1$ the accretion process by primordial black holes is robust, even if they are highly spinning initially.
The second scenario corresponds to a supermassive black hole which accretes dark matter from its environment. Assuming a reservoir of constant density and low temperature and integrating Eq. (23), one finds that the time required for the black hole mass to grow by a factor $X$ (and the black hole spin parameter to decrease from $\alpha_{0}$ to $\alpha=\alpha_{0}/X^{2+2\delta}$) is $t_{X}=(1-1/X)c^{5}/(G^{2}M_{0}\mathcal{A}\varepsilon_{\infty})$, where for simplicity we have set $t_{0}=0$. Therefore, the characteristic time scale $t_{\infty}$ for having a significant mass growth is related to the parameter $z$ according to
$\sqrt{z}=\frac{c^{5}}{\sqrt{512\pi}G^{2}M_{0}\varepsilon_{\infty}t_{\infty}},$ (25)